OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6). Answer 0 14 7 -14

OpenStudy (anonymous):

To calculate the slope of a line given two points, you use the rise/run formula, which is: y2 - y1 / x2 - x1 Then you'll have your answer.

OpenStudy (anonymous):

i know that, but i don't know how to do it.

OpenStudy (anonymous):

That means -6 - 8 which is -14 Divided by -3 - -4 which is the same as -3 +4 = 1 -14/1 is -14

OpenStudy (anonymous):

When two points are given in question like (5,6) (2,3) so when we are viewing the formula m=y2-y1/x2-x1 it means that we have to plug the values of our points in this formula: In my given example 3=y2, 6=y1 / 2=x2, 5=x1 Now when I plug these values it will look something like this: m =3-6/2-5 which will be equal to -3/-3 = 1

OpenStudy (anonymous):

Out Slope in my given example = 1

OpenStudy (anonymous):

oh okay , this is just hard for me to understand , but thanks .

OpenStudy (anonymous):

when given two points in question, see those points like this (x1,y1) (x2,y2) and you know the formula of finding the slope: m= y2-y1 / x2-x1 Just plug the points. It is really simple idea. If you are still not understanding this, I would say watch this video by Sal Khan .. http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/v/slope-of-a-line

OpenStudy (anonymous):

Yeah its just that one equation you have to remember "rise over run" meaning find the difference in the y-axes divided by the x axes. Although this does assume that you know how to add and subtract negative numbers in combination with positive numbers.

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

-14 if u still need it lol